# On the electron self-energy to three loops in QED

### v1

@@ Abstract
We compute the electron self-energy in Quantum Electrodynamics to three loops in terms of iterated integrals over kernels of elliptic type. We make use of the differential equations method, augmented by an $\epsilon$-factorized basis, which allows us to gain full control over the differential forms appearing in the iterated integrals to all orders in the dimensional regulator. We obtain compact analytic expressions, for which we provide generalized series expansion representations that allow us to evaluate the result numerically for all values of the electron momentum squared. As a by product, we also obtain $\epsilon$-resummed results for the self-energy in the on-shell limit $p^{2}=m^{2}$, which we use to recompute the known three-loop renormalization constants in the on-shell scheme.
@@ 1 Introduction
From the first calculations by Schwinger and Feynman, [^1] See ref. Bovy:2006dr for a summary of these historical developments. a lot has changed in the way we address the calculation of loop Feynman integrals. In fact, the increasing precision of high- and low-energy particle experiments on the one hand, and on the other the formal interest in searching for general mathematical patterns in perturbative QFT, have triggered the development of new mathematical methods which have enormously simplified these calculations. These include integration-by-parts identities Tkachov:1981wb ; Chetyrkin:1981qh and the Laporta algorithm Laporta:2004rb , the differential equation method Kotikov:1990kg ; Remiddi:1997ny ; Gehrmann:1999as , and a new understanding of the theory of the special functions involved, in particular multiple polylogarithms Kummer ; Remiddi:1999ew ; Goncharov:1995 ; Goncharov:1998kja ; Vollinga:2004sn ; Duhr:2011zq , and more general iterated integrals ChenSymbol over logarithmic differential forms (dlog forms). As a result, extremely sophisticated calculations have become possible, including scattering amplitudes for processes involving three, four or five external particles, up to four, three and two loops, respectively, see refs. FebresCordero:2022psq ; Caola:2022ayt and references therein.
@@
Despite these impressive results, up until recently a fully analytic result for the two-loop corrections to the very fundamental (and apparently simpler) electron self-energy were not available. The first attempt to this calculation can be traced back to A. Sabry Sabry in 1962, who showed that the discontinuity of the self-energy contains new mathematical objects which cannot be expressed as iterated integrals over dlog forms, namely elliptic integrals. While Sabry could provide series representations for the two-loop self-energy, we had to wait more than half a century for its first fully analytic calculation Honemann:2018mrb . This became possible thanks to recent advances in our understanding of iterated integrals defined over more general classes of differential forms, in particular those defined over genus-one curves. If the genus-one curve has in addition one marked point, the curve is also known as an elliptic curve. Interestingly, while these new geometries can give rise to extremely rich mathematical structures, the two-loop electron self-energy turned out to be expressible in terms of a very well-defined subclass of these integrals, whose differential forms have especially simple transformation properties under a subgroup of modular transformations, namely iterated integrals of modular forms for the congruence subgroup $\Gamma_{1}(6)$ Adams:2017ejb .
@@
In turn, standard iterated integrals over modular forms are well known not to be sufficient to describe elliptic scattering amplitudes with non-trivial kinematic dependence already at two loops, and it remains an open question if a well-defined extension of these integrals can accommodate a large enough number of physically relevant problems. An especially powerful technique to investigate the space of functions characterizing a given physical problem is the differential equations method, in particular when augmented by the idea of a canonical basis ArkaniHamed:2010gh ; Henn:2013pwa . These are bases of master integrals which fulfill especially simple differential equations, where the dependence on the dimensional regulator $\epsilon$ is entirely factorized from the dependence on the kinematics. In addition, if the differential forms appearing in the connection matrix are dlog forms, the corresponding solutions can be typically expressed as iterated integrals of polylogarithmic type. [^2] Though care is needed, because there may be exceptions, where an iterated integral of dlog-forms cannot be expressed in terms of polylogarithms Duhr:2020gdd . While the notion of a canonical basis remained for long time confined to dlog forms, recently important steps forward have been made to understand its generalization to elliptic and Calabi-Yau geometries Adams:2016xah ; Adams:2018bsn ; Pogel:2022ken ; Frellesvig:2021hkr ; Dlapa:2022wdu ; Frellesvig:2023iwr ; Gorges:2023zgv ; Pogel:2022yat . A basis in this form allows one not only to make general considerations about the mathematical structure of the solutions to all orders in the dimensional regulator, but also provides a way to investigate the analytic properties of the solution close to the physical number of space-time dimensions $D=4$.
@@ 2 Definitions and Feynman diagrams calculation
It turns out that, after using symmetry identities and integration-by-parts (IBP) identities, family $B$ does not introduce any additional master integral that cannot be expressed in terms of family $A$. We therefore ignore it in what follows. We reduce all scalar integrals to master integrals using [^3] In part of the calculation also the package LiteRed Lee:2012cn was used. Reduze 2 Studerus:2009ye ; vonManteuffel:2012np and Kira 2 Maierhofer:2017gsa ; Klappert:2019emp ; Klappert:2020nbg using IBP identities and we find a $51$-dimensional basis, as quoted above, all from family $A$:
@@ 2.1 The elliptic curve
In this specific case, there are four regular singular points $x\in\{0,1,9,\infty\}$. For convenience, we provide here our choice of the solutions close to each singular point, in terms of their series expansion through the first four orders.
@@ 3 Differential equations in $\epsilon$-form
where $\mathbb{P}$ is the path-ordering operator, $\vec{J}_{0}$ is the boundary condition at $x=x_{0}$ and $\gamma$ is a path that connects the points $x_{0}$ to the generic point $x$. In this form, it becomes obvious that at all orders in $\epsilon$ the integrals $\vec{J}$ can be written as linear combinations of (Chen) iterated integrals ChenSymbol over the forms $\omega_{i}$. In the problem under study we only deal with one kinematic variable and the ensuing iterated integrals are explicitly given by
$I(\omega_{i_{n}},\ldots,\omega_{i_{1}})=\int_{x_{0}}^{x}\mathrm{d}x_{n}f_{i_{n}}(x_{n})\ldots\int_{x_{0}}^{x_{3}}\mathrm{d}x_{2}f_{i_{2}}(x_{2})\int_{x_{0}}^{x_{2}}\mathrm{d}x_{1}f_{i_{1}}(x_{1})\,,$ (30)
where $x_{0}$ is the chosen boundary point.
@@
This is the statement that the $\omega_{i}$’s form a basis of a vector space. Equation (37) can alternatively be proved directly by comparing the local series expansion representations. We notice that these relations do not exclude more subtle relations, where linear combinations of the differential forms could be expressed as a total differential of an arbitrarily complicated function. Nevertheless, we believe that the conditions in eq. (36) should be enough to guarantee that no combination of the iterated integrals originating from these forms evaluates trivially to zero or can be reduced to a lower length iterated integral over the same set of forms.
@@ 4 Solution in terms of iterated integrals
While we could recover most boundary values close to $x=1$ from the analytic results for the master integrals provided in refs. Laporta:1996mq ; Melnikov:1999xp , we decided to follow a different approach that can be more easily generalized to other boundary points as well. In order to obtain the limiting value of the master integrals close to a regular singular point $x=x_{0}$, we use AMFlow Liu:2022chg to obtain a high precision evaluation close to that point. For our problem, we typically choose $x=x_{0}\pm\delta$ with $\delta\sim\mathcal{O}(10^{-2})$. We can then use the differential equations to transport this boundary value back to exactly $x=x_{0}$. Let us assume for definiteness that we perform the matching to the right of the singular point. Then we achieve this by computing, order by order in $\epsilon$, a generalized series expansion solution for the master integrals close to $x=x_{0}^{+}$ with an arbitrary number of terms and matching this to the numerical result obtained at $x=x_{0}+\delta$. In formulas, if we expand the master integrals in $\epsilon$ as
@@ 4.1 General structure of the solution in terms of iterated integrals
$\Sigma_{V,-1}^{(3)}$ $=R_{1}+R_{2}I(f_{4})+\frac{R_{3}}{\varpi_{0}}I(f_{11})+R_{4}I(f_{4},f_{4})+R_{5}\Big{[}6I(f_{3},f_{4})-\pi^{2}\Big{]}$
$+\left(R_{6}\varpi_{0}+R_{7}\varpi_{0}^{\prime}\right)\left[9I(f_{10},f_{11})-2\sqrt{3}\text{Cl}_{2}(\tfrac{\pi}{3})\right]+R_{8}\Big{[}144I(f_{12},f_{10},f_{11})$
$+36I(f_{3},f_{4},f_{4})-18I(f_{11},f_{10},f_{11})-18I(f_{4},f_{3},f_{4})+3\pi^{2}I(f_{4})$
$+4\sqrt{3}\text{Cl}_{2}(\tfrac{\pi}{3})\big{(}I(f_{11})-8I(f_{12})\big{)}\Big{]}\,,$
@@
$\Sigma_{S,-2}^{(3)}$ $=-\frac{673+56x}{864(1-x)}-\frac{437-5x}{384x}I(f_{4})\,,$
$\Sigma_{S,-1}^{(3)}$ $=\tilde{R}_{1}+\tilde{R}_{2}\Big{[}7\pi^{2}-144\zeta(3)\Big{]}+\tilde{R}_{3}I(f_{4})+\frac{\tilde{R}_{4}}{\varpi_{0}}I(f_{11})+\tilde{R}_{5}I(f_{4},f_{4})\,,$
$+\tilde{R}_{6}\Big{[}7I(f_{3},f_{4})-24\zeta(3)\Big{]}+\left(\tilde{R}_{7}\varpi_{0}+\tilde{R}_{8}\varpi_{0}^{\prime}\right)\Big{[}9I(f_{10},f_{11})-2\sqrt{3}\text{Cl}_{2}(\tfrac{\pi}{3})\Big{]}$
$+\tilde{R}_{9}\Big{[}1008I(f_{12},f_{10},f_{11})+252I(f_{3},f_{4},f_{4})-126I(f_{11},f_{10},f_{11})-126I(f_{4},f_{3},f_{4})$
$+28\sqrt{3}\,\text{Cl}_{2}(\tfrac{\pi}{3})\big{(}I(f_{11})-8I(f_{12})\big{)}+21\pi^{2}I(f_{4})-2484\zeta(3)\Big{]}\,,$ (48)
@@ 4.2 Local solution close to $p^{2}=m^{2}$
The behavior of the self-energy for $p\to m^{2}$ is physically relevant, as it is central to determine the mass and wave function renormalization constants in the on-shell scheme. As it is well known, $p^{2}$ acts in this case as an infrared regulator, and taking the on-shell limit produces new infrared divergences, which render the extraction of the renormalization constants from the full result non-trivial. As discussed above, see eq. (43), from our $\epsilon$-factorized differential equations, it is easy to obtain a fully $\epsilon$-resummed expression for the self-energy coefficients close to $p^{2}=m^{2}$. This makes the branch structure of the self-energy manifest and also allows us to take the on-shell limit at once, simply by discarding the divergent branches.
@@ 4.3 Local solution close to $p^{2}=0$
Let us make some comment about the appearance of $\sqrt{3}$ starting from two loops. At least up to three loops, the self-energy at $p^{2}=0$ can be expressed in terms of multiple polylogarithms evaluated at sixth roots of unity, i.e., multiple polylogarithms where all arguments are either 0 or $\rho^{k}$, with $60\leq k<$ 6 and $\rho=e^{i\pi/3}=\tfrac{1}{2}+\tfrac{i\sqrt{3}}{2}$. Consequently, it is natural to separate the transcendental numbers into their real and imaginary parts. This defines a natural notion of parity on these numbers, under which the real parts are even, and the imaginary parts are odd. If we define $\sqrt{3}=2\,\textrm{Im}(\rho)$ to be odd under this parity (and rational numbers are even), we see that the self-energy is even, as expected.
@@ 6 Conclusions
The series expansion representation for the solution has instead been used to obtain fast and reliable numerics for the self-energy across the whole parameter space. While we have typically produced series representation using numerical boundary conditions in the various regions of interest, we have studied two points in more detail, namely $p^{2}=0$ and $p^{2}=m^{2}$. For what concerns $p^{2}=0$, we have shown that the self-energy can be expressed in terms of multiple polylogarithms evaluated at sixth roots of unity. We have provided an alternative basis for these numbers up to weight four, which allows us to maximally express all constants in terms of classical polylogarithms. We have then dedicated special attention to derive an $\epsilon$-resummed series representation close to the point $p^{2}=m^{2}$, which allowed us to obtain at once both the mass and the wave function renormalization constants in the on-shell scheme. Here we have shown that, as expected, our results up to finite order in $\epsilon$ can be written in terms of multiple zeta values and classical polylogarithms evaluated at $x=\frac{1}{2}$. Finally, we have studied the effect of introducing Bernoulli-like variables on the convergence of the series expansion and observed that, while they substantially accelerate the convergence of the series expansions at most regular singular points, they do not seem to work equally well everywhere, and instead they appear to deteriorate the convergence of the series close to the on-shell limit $p^{2}=m^{2}$.

### v2

@@ Abstract
We compute the electron self-energy in Quantum Electrodynamics to three loops in terms of iterated integrals over kernels of elliptic type. We make use of the differential equations method, augmented by an $\epsilon$-factorized basis, which allows us to gain full control over the differential forms appearing in the iterated integrals to all orders in the dimensional regulator. We obtain compact analytic expressions, for which we provide generalized series expansion representations that allow us to evaluate the result numerically for all values of the electron momentum squared. As a by-product, we also obtain $\epsilon$-resummed results for the self-energy in the on-shell limit $p^{2}=m^{2}$, which we use to recompute the known three-loop renormalization constants in the on-shell scheme.
@@ 1 Introduction
From the first calculations by Schwinger and Feynman (see ref. Bovy:2006dr for a summary of these historical developments), a lot has changed in the way we address the calculation of loop Feynman integrals. In fact, the increasing precision of high- and low-energy particle experiments on the one hand, and on the other the formal interest in searching for general mathematical patterns in perturbative QFT, have triggered the development of new mathematical methods which have enormously simplified these calculations. These include integration-by-parts identities Tkachov:1981wb ; Chetyrkin:1981qh and the Laporta algorithm Laporta:2004rb , the differential equation method Kotikov:1990kg ; Remiddi:1997ny ; Gehrmann:1999as , and a new understanding of the theory of the special functions involved, in particular multiple polylogarithms Kummer ; Remiddi:1999ew ; Goncharov:1995 ; Goncharov:1998kja ; Vollinga:2004sn ; Duhr:2011zq , and more general iterated integrals ChenSymbol over logarithmic differential forms (dlog forms). As a result, extremely sophisticated calculations have become possible, including scattering amplitudes for processes involving three, four or five external particles, up to four, three and two loops, respectively, see refs. FebresCordero:2022psq ; Caola:2022ayt and references therein.
@@
Despite these impressive results, up until recently a fully analytic result for the two-loop corrections to the very fundamental (and apparently simpler) electron self-energy were not available. The first attempt to this calculation can be traced back to A. Sabry Sabry in 1962, who showed that the discontinuity of the self-energy contains new mathematical objects which cannot be expressed as iterated integrals over dlog forms, namely elliptic integrals. While Sabry could provide series representations for the two-loop self-energy, we had to wait more than half a century for its first fully analytic calculation Honemann:2018mrb . This became possible thanks to recent advances in our understanding of iterated integrals defined over more general classes of differential forms, in particular those defined over genus-one curves. If the genus-one curve has in addition one marked point, the curve is also known as an elliptic curve MR2024529 . Interestingly, while these new geometries can give rise to extremely rich mathematical structures, the two-loop electron self-energy turned out to be expressible in terms of a very well-defined subclass of these integrals, whose differential forms have especially simple transformation properties under a subgroup of modular transformations, namely iterated integrals of modular forms for the congruence subgroup $\Gamma_{1}(6)$ Adams:2017ejb .
@@
In turn, standard iterated integrals over modular forms are well known not to be sufficient to describe elliptic scattering amplitudes with non-trivial kinematic dependence already at two loops, and it remains an open question if a well-defined extension of these integrals can accommodate a large enough number of physically relevant problems. An especially powerful technique to investigate the space of functions characterizing a given physical problem is the differential equation’s method, in particular when augmented by the idea of a canonical basis ArkaniHamed:2010gh ; Henn:2013pwa . These are bases of master integrals which fulfill especially simple differential equations, where the dependence on the dimensional regulator $\epsilon$ is entirely factorized from the dependence on the kinematics. In addition, if the differential forms appearing in the connection matrix are dlog forms, the corresponding solutions can be typically expressed as iterated integrals of polylogarithmic type. Though care is needed, because there may be exceptions, where an iterated integral of dlog forms cannot be expressed in terms of polylogarithms Duhr:2020gdd . While the notion of a canonical basis remained for long time confined to dlog forms, recently important steps forward have been made to understand its generalization to elliptic and Calabi-Yau geometries Adams:2016xah ; Adams:2018bsn ; Pogel:2022ken ; Frellesvig:2021hkr ; Dlapa:2022wdu ; Frellesvig:2023iwr ; Gorges:2023zgv ; Pogel:2022yat ; Behring:2023rlq . A basis in this form allows one not only to make general considerations about the mathematical structure of the solutions to all orders in the dimensional regulator, but also provides a way to investigate the analytic properties of the solution close to the physical number of space-time dimensions $D=4$.
@@ 2 Definitions and Feynman diagrams calculation
It turns out that, after using symmetry identities and integration-by-parts (IBP) identities, family $B$ does not introduce any additional master integral that cannot be expressed in terms of family $A$. We therefore ignore it in what follows. We reduce all scalar integrals to master integrals using Reduze 2 Studerus:2009ye ; vonManteuffel:2012np , Kira 2 Maierhofer:2017gsa ; Klappert:2019emp ; Klappert:2020nbg and also partly LiteRed Lee:2012cn using IBP identities and we find a $51$-dimensional basis, as quoted above, all from family $A$:
@@ 2.1 The elliptic curve
In this specific case, there are four regular singular points $x\in\{0,1,9,\infty\}$. For convenience, we provide here our choice of the solutions close to each singular point, in terms of their series expansion through the first four orders. Our basis is chosen such that the branch cut of the logarithmic solution is on the negative real axis of the local variable.
@@ 3 Differential equations in $\epsilon$-form
where $\mathbb{P}$ is the path-ordering operator, $\vec{J}_{0}$ is the boundary condition at $x=x_{0}$ and $\gamma$ is a path that connects the points $x_{0}$ to the generic point $x$. In this form, it becomes obvious that at all orders in $\epsilon$ the integrals $\vec{J}$ can be written as linear combinations of (Chen) iterated integrals ChenSymbol over the forms $\omega_{i}$. In the problem under study we only deal with one kinematic variable and the ensuing iterated integrals are given by
$I(f_{i_{n}},\ldots,f_{i_{1}};x)=\int_{0}^{x}\mathrm{d}x_{n}\,f_{i_{n}}(x_{n})\ldots\int_{0}^{x_{3}}\mathrm{d}x_{2}\,f_{i_{2}}(x_{2})\int_{0}^{x_{2}}\mathrm{d}x_{1}\,f_{i_{1}}(x_{1})\,,$ (30)
where we fixed the boundary point at $x_{0}=0$ and we assume that shuffle regularization is used throughout to make sense of the divergent differential forms.
@@
This is the statement that the $\omega_{i}$’s form a basis of a vector space. Equation (37) as well as (36) can alternatively be proved directly by comparing the local series expansion representations. Up to 200 orders in the expansion of the differential forms $\omega_{i}$ and Frobenius solutions $\varpi_{0}$ and $\partial_{x}\varpi_{0}$, we have not found any relation of these types. We notice that these relations do not exclude more subtle relations, where linear combinations of the differential forms could be expressed as a total differential of an arbitrarily complicated function. Nevertheless, we believe that the conditions in eq. (36) should be enough to guarantee that no combination of the iterated integrals originating from these forms evaluates trivially to zero or can be reduced to a lower length iterated integral over the same set of forms.
@@ 4 Solution in terms of iterated integrals
While we could recover most boundary values close to $x=1$ from the analytic results for the master integrals provided in refs. Laporta:1996mq ; Melnikov:1999xp , we decided to follow a different approach that can be more easily generalized to other boundary points as well. In order to obtain the limiting value of the master integrals close to a regular singular point $x=x_{0}$, we use AMFlow Liu:2017jxz ; Liu:2022chg to obtain a high precision evaluation close to that point. For our problem, we typically choose $x=x_{0}\pm\delta$ with $\delta\sim\mathcal{O}(10^{-2})$. We can then use the differential equations to transport this boundary value back to exactly $x=x_{0}$. Let us assume for definiteness that we perform the matching to the right of the singular point. Then we achieve this by computing, order by order in $\epsilon$, a generalized series expansion solution for the master integrals close to $x=x_{0}^{+}$ with an arbitrary number of terms and matching this to the numerical result obtained at $x=x_{0}+\delta$. In formulas, if we expand the master integrals in $\epsilon$ as
@@ 4.1 General structure of the solution in terms of iterated integrals
$\Sigma_{V,-1}^{(3)}$ $=R_{1}+R_{2}I(f_{4};x)+\frac{R_{3}}{\varpi_{0}}I(f_{11};x)+R_{4}I(f_{4},f_{4};x)+R_{5}\Big{[}6I(f_{3},f_{4};x)-\pi^{2}\Big{]}$
$+\left(R_{6}\varpi_{0}+R_{7}\varpi_{0}^{\prime}\right)\left[9I(f_{10},f_{11};x)-2\sqrt{3}\text{Cl}_{2}(\tfrac{\pi}{3})\right]+R_{8}\Big{[}144I(f_{12},f_{10},f_{11};x)$
$+36I(f_{3},f_{4},f_{4};x)-18I(f_{11},f_{10},f_{11};x)-18I(f_{4},f_{3},f_{4};x)+3\pi^{2}I(f_{4};x)$
$+4\sqrt{3}\text{Cl}_{2}(\tfrac{\pi}{3})\big{(}I(f_{11};x)-8I(f_{12};x)\big{)}\Big{]}\,,$
@@
$\Sigma_{S,-2}^{(3)}$ $=-\frac{673+56x}{864(1-x)}-\frac{437-5x}{384x}I(f_{4};x)\,,$
$\Sigma_{S,-1}^{(3)}$ $=\tilde{R}_{1}+\tilde{R}_{2}\Big{[}7\pi^{2}-144\zeta(3)\Big{]}+\tilde{R}_{3}I(f_{4};x)+\frac{\tilde{R}_{4}}{\varpi_{0}}I(f_{11};x)+\tilde{R}_{5}I(f_{4},f_{4};x)\,,$
$+\tilde{R}_{6}\Big{[}7I(f_{3},f_{4};x)-24\zeta(3)\Big{]}+\left(\tilde{R}_{7}\varpi_{0}+\tilde{R}_{8}\varpi_{0}^{\prime}\right)\Big{[}9I(f_{10},f_{11};x)-2\sqrt{3}\text{Cl}_{2}(\tfrac{\pi}{3})\Big{]}$
$+\tilde{R}_{9}\Big{[}1008I(f_{12},f_{10},f_{11};x)+252I(f_{3},f_{4},f_{4};x)-126I(f_{11},f_{10},f_{11};x)$
$-126I(f_{4},f_{3},f_{4};x)+28\sqrt{3}\,\text{Cl}_{2}(\tfrac{\pi}{3})\big{(}I(f_{11};x)-8I(f_{12};x)\big{)}$
$+21\pi^{2}I(f_{4};x)-2484\zeta(3)\Big{]}\,,$ (48)
@@ 4.2 Local solution close to $p^{2}=m^{2}$
The behavior of the self-energy for $p^{2}\to m^{2}$ is physically relevant, as it is central to determine the mass and wave function renormalization constants in the on-shell scheme. As it is well known, $p^{2}$ acts in this case as an infrared regulator, and taking the on-shell limit produces new infrared divergences, which render the extraction of the renormalization constants from the full result non-trivial. As discussed above, see eq. (43), from our $\epsilon$-factorized differential equations, it is easy to obtain a fully $\epsilon$-resummed expression for the self-energy coefficients close to $p^{2}=m^{2}$. This makes the branch structure of the self-energy manifest and also allows us to take the on-shell limit at once, simply by discarding the divergent branches.
@@ 4.3 Local solution close to $p^{2}=0$
Let us make some comment about the appearance of $\sqrt{3}$ starting from two loops. At least up to three loops, the self-energy at $p^{2}=0$ can be expressed in terms of multiple polylogarithms evaluated at sixth roots of unity, i.e., multiple polylogarithms where all arguments are either 0 or $\rho^{k}$, with $60\leq k<$ 6 and $\rho=e^{i\pi/3}=\tfrac{1}{2}+\tfrac{i\sqrt{3}}{2}$. Consequently, it is natural to separate the transcendental numbers into their real and imaginary parts. This defines a natural notion of parity on these numbers, under which the real parts are even, and the imaginary parts are odd. If we define $\sqrt{3}=2\,\textrm{Im}(\rho)$ to be odd under this parity (and rational numbers are even), we see that the self-energy is even. The fact that the self-energy is parity-even can be seen as follows: The Feynman integrals that define the self-energy at a given loop order are integrals of rational functions, and thus independent of the choice of the sign of any square root that may appear in the answer. In particular, it cannot depend on the choice of sign for the square root of 3, and thus it is parity-even.
@@ 6 Conclusions
The series expansion representation for the solution has instead been used to obtain fast and reliable numerics for the self-energy across the whole parameter space. While we have typically produced series representation using numerical boundary conditions in the various regions of interest, we have studied two points in more detail, namely $p^{2}=0$ and $p^{2}=m^{2}$. For what concerns $p^{2}=0$, we have shown that the self-energy can be expressed in terms of multiple polylogarithms evaluated at sixth roots of unity. We have provided an alternative basis for these numbers up to weight four, which allows us to maximally express all constants in terms of classical polylogarithms. We have then dedicated special attention to derive an $\epsilon$-resummed series representation close to the point $p^{2}=m^{2}$, which allowed us to obtain at once both the mass and the wave function renormalization constants in the on-shell scheme. Here we have shown that, as expected, our results up to finite order in $\epsilon$ can be written in terms of multiple zeta values and classical polylogarithms evaluated at $x=\frac{1}{2}$. Finally, we have studied the effect of introducing Bernoulli-like variables on the convergence of the series expansion and observed that, while they substantially accelerate the convergence of the series expansions at most regular singular points, they do not seem to work equally well everywhere. The least improvement is observed close to the on-shell limit $p^{2}=m^{2}$.